What value for y will make the equation true?
4(y-2)+6y = 2(3y +2)

Answer by fractalier(6550)

We solve this equation
4(y-2)+6y = 2(3y +2)
Expand and collect like terms first...
4y - 8 + 6y = 6y + 4
10y - 8 = 6y + 4
Now subtract 6y and add 8 and get
4y = 12
y = 3